# MonoForce: Learnable Image-conditioned Physics Engine

### v3

@@ III-B Differentiable Physics Engine
When the robot is moving over a terrain, two types of external forces are acting on the point cloud $\mathcal{P}$ representing its model: (i) gravitational forces and (ii) robot-terrain interaction forces. The former is defined as $m_{i}\mathbf{g}=[0,0,-m_{i}g]^{\top}$ and acts on all the points of the robot at all times, while the latter is the result of complex physical interactions that are not easy to model explicitly and act only on the points of the robot that are in contact with the terrain. There are two types of robot-terrain interaction forces: (i) normal terrain force that prevents the penetration of the terrain by the robot points, (ii) tangential friction force that generates forward acceleration when the tracks are moving, and prevents side slippage of the robot.
Robot-terrain interaction forces
@@
Normal reaction forces.
One extreme option is to predict the 3D force vectors $\mathbf{F}_{i}$ directly by a neural network, but we decided to enforce additional prior assumptions to reduce the risk of overfitting. These prior assumptions are based on common intuition from the contact dynamics of flexible objects. In particular, we assume that the magnitude of the force that the terrain exerts on the point $\mathbf{p}_{i}\in\mathcal{P}$ increases proportionally to the deformation of the terrain. Consequently, the network does not directly predict the force, but rather predicts the height of the terrain $h\in\mathcal{H}_{t}$ at which the force begins to act on the robot body and the stiffness of the terrain $e\in\mathcal{K}$. We understand the quantity $e$ as an equivalent of the spring constant from Hooke’s spring model, Figure 8. Given the stiffness of the terrain and the point of the robot that penetrated the terrain by ${\Delta}h$, the reaction force is calculated as $e\cdot{\Delta}h$.
@@
Tangential friction forces.
Our tracked robot navigates by moving the main tracks and 4 flippers (auxiliary tracks). The flipper motion is purely kinematic in our model. This means that in a given time instant, their pose is uniquely determined by a $4$-dimensional vector of their rotations, and they are treated as a rigid part of the robot. The motion of the main tracks is transformed into forces tangential to the terrain. The friction force delivers forward acceleration of the robot when robot tracks (either on flippers or on main tracks) are moving. At the same time, it prevents the robot from sliding sideways. When a robot point $\mathbf{p}_{i}$, which belongs to a track, is in contact with terrain with friction coefficient $\mu\in\mathcal{M}$, the resulting friction force at a contact point is computed as follows, [33]:
@@
Implementation of the Differentiable ODE Solver
We implement the robot-terrain interaction ODE (3) in PyTorch [43]. The Neural ODE framework [44] is used to solve the system of ODEs. For efficiency reasons, we utilize the Euler integrator for the ODE integration. The differentiable ODE solver [44] estimates the gradient through the implicit function theorem. Additionally, we implement the ODE (3) solver that estimates gradient through auto-differentiation [43], i.e. it builds and retains the full computational graph of the feedforward integration.
@@ III-C Data-driven Trajectory Prediction
Given an initial robot’s state $\mathbf{x}_{0}$ and a sequence of control inputs for a time horizon $T$, $\mathbf{u}_{t},t\in\{0\dots T\}$, the TrajLSTM model provides a sequence of states at control command time moments, $\mathbf{x}_{t},t\in\{0\dots T\}$. As in outdoor scenarios the robot commonly traverses uneven terrain, we additionally include the terrain shape input to the model in the form of heightmap $\mathcal{H}=\mathcal{H}_{0}$ estimated at initial time moment $t=0$. Each timestep’s control input $\mathbf{u}_{i}$ is concatenated with the shared spatial features $\mathbf{x}_{i}$, as shown in Figure 10. The combined features are passed through dense layers to prepare for temporal processing. The LSTM unit [28] processes the sequence of features (one for each timestep). As in our experiments, the time horizon for trajectory prediction is reasonably small, $T=5[$\sec$]$, and the robot’s trajectories lie within the heightmap area, we use the shared heightmap input for all the LSTM units of the network. So the heightmap is processed through the convolutional layers once and flattened, producing a fixed-size spatial feature vector. This design choice (of not processing the heightmaps at different time moments) is also motivated by computational efficiency reason. At each moment $t$, this heightmap vector is concatenated with the fused spatial-control features and processed by an LSTM unit. The LSTM unit output for each timestep $t$ is passed through a fully connected (dense) layer to produce the next state $\mathbf{x}_{t+1}$. The sequence of states form the predicted trajectory, $\{\mathbf{x}_{0},\dots\mathbf{x}_{T}\}$.
@@ III-D End-to-end Learning
Since the architecture Figure 2 is end-to-end differentiable, we can directly learn to predict all intermediate outputs just using trajectory loss (8). An example of terrain learning with the trajectory loss is visualized in Figure 9. To make the training more efficient and the learned model explainable, we employ the geometrical loss (9) and terrain loss (10) as regularization terms. stat
Figure 11: MonoForce prediction examples. Left: The robot is moving through a narrow passage between a wall and tree logs. Right: The robot is moving on a gravel road with rocks on the sides. It starts its motion from the position marked with a coordinate frame and the trajectory is predicted for $10~{}[$\sec$]$ using real control commands. The camera images are taken from the robot’s initial position (top row). The visualization includes predicted supporting terrain $\mathcal{H}_{t}$ (second row). It is additionally shown in 3D and colored with predicted friction values (third row).
@@
The Figure 11 show the prediction examples of the MonoForce model in diverse outdoor environments. From the example on the left, we can see that the model correctly predicts the robot’s trajectory and the terrain shape suppressing traversable vegetation, while the rigid obstacles (wall and tree logs) are correctly detected. The example on the right demonstrates the model’s ability to predict the robot’s trajectory ($10~{}[$\sec$]$-long horizon) with reasonable accuracy and to detect the rigid obstacles (stones) on the terrain. It could also be noticed that the surfaces that provide the robot good traction (paved and gravel roads) are marked with a higher friction value, while for the objects that might not give good contact with the robot’s tracks (walls and tree logs) the friction value is lower.
@@
We argue that the friction estimates are approximate and an interesting research direction could be comparing them with real-world measurements or with the values provided by a high-fidelity physics engine (e.g. AGX Dynamics [37]). However, one of the benefits of our differentiable approach is that the model does not require ground-truth friction values for training. The predicted heightmap’s size is $12.8\times 12.8$\mathrm{m}$^{2}$ and the grid resolution is $0.1$\mathrm{m}$$. It has an upper bound of $1~{}[$\mathrm{m}$]$ and a lower bound of $-1~{}[$\mathrm{m}$]$. This constraint was introduced based on the robot’s size and taking into account hanging objects (tree branches) that should not be considered as obstacles (Figure 17). Additionally, the terrain is predicted in the gravity-aligned frame. That is made possible thanks to the inclusion of camera intrinsics and extrinsics as input to the model, Figure 6. It also allows correctly modeling the robot-terrain interaction forces (and thus modeling the robot’s trajectory accurately) for the scenarios with non-flat terrain, for example, going uphill or downhill. This will not be possible if only camera images are used as input.
IV Experiments and Results
IV-A Dataset: ROUGH
@@
Despite the amount and annotation quality of the data provided in the RELLIS-3D sequences, it lacks examples of a robot moving over hills, obstacles, and traversing high grass. To fill this gap in autonomous off-road navigation, a new Rugged Obstacles and Uneven Ground Hardcore (ROUGH) dataset will be released, containing sensory data from forest and field scenarios recorded with the two robot platforms: MARV, Bluebotics AbsolemFigure 12. It contains several hours of driving with the mid-sized robots in challenging terrain. The ultimate goal is to provide recordings of traversals on (or through) flexible natural obstacles like tall grass, hanging tree branches, mud, dense undergrowth, etc. The ROUGH sequences are collected with shape-changing tracked robots (Figure 12), which allows capturing a much larger range of dynamic responses - simply moving the auxiliary tracks changes the center and moments of inertia and the contact surface. To correctly utilize the effects of dynamics, ROUGH provides not only sensory data but also robot models with dynamic properties. The data set also contains carefully time-synchronized point cloud scans from Ouster OS0-128 lidar and corresponding RGB images from $4$ Basler cameras installed on the robots (front, rear, left, and right). The robot localization data obtained using ICP SLAM [47] is also recorded. It was a better choice than GPS which is unreliable near or under tree canopy and close to buildings. The controls ($\mathbf{u}$ in Figure 6) in the form of commanded tracks velocities and angles are recorded and used to train and evaluate the MonoForce model.
After training the MonoForce terrain encoder on the RELLIS-3D dataset, we fine-tune it on the ROUGH data sequences. Please note that the dataset is under construction at the moment of writing and will be extended with the help of new robot platforms and data collection scenarios. The ROUGH dataset will be made publicly available [^2] https://github.com/ctu-vras/rough-dataset.
@@ IV-B Sensor Fusion
Figure 13: BEVFusion [22] architecture. The model processes multi-modal inputs (RGB images and lidar point cloud) and fuses them into a shared bird’s-eye view (BEV) space. The LSS [20] architecture part is sketched in green (upper branch), while the model branch processing point cloud input (VoxelNet [21]) is depicted in blue. The fused BEV features are further used for terrain properties prediction tasks.
In this section, we study to what extent the inclusion of other sensor modalities (except for RGB images) improves terrain estimation accuracy. As we discuss in Section III-A, the terrain encoder model predicts terrain properties solely from RGB input. However, a natural question arises whether an additional point cloud input or a sensor fusion technique could improve the prediction. We adapt the architecture introduced in [22] for terrain properties prediction. The modified BEVFusion architecture for the terrain prediction task from RGB and lidar inputs is provided in the Figure 13.
For different sensor inputs (multi-view RGB cameras and lidar), modality-specific encoders are applied to extract their features. Then the sensor-specific features are transformed into a unified BEV representation that preserves both semantic and geometric information. Following the [22] approach the convolution-based BEV encoder is applied next to the stacked BEV features to account for possible local misalignment between different sensors. The task-specific heads are added to predict different terrain properties (for example, its shape as a form of heightmap and friction) as shown in the Figure 13. The LSS [20] model is used as the camera input BEV-encoder, as described in Section III-A and the BEVFusion [22] paper. The VoxelNet [21] model is utilized as the point cloud input BEV-encoder. First, the raw point cloud input is voxelized into a $3D$ voxel grid. The point cloud coordinates are mapped to a discrete grid of binary indices (a value of $1$ denotes that the voxel is occupied). Next, the $3D$-convolution-based lidar Encoder processes the voxelized input to extract high-level feature representation. This operation is followed by $\max$-pooling along the $z$-axis to reduce the $3D$ feature map to $2D$ BEV space.
We train and evaluate the LSS [20], VoxelNet [21], and BEVFusion [22] terrain encoders on the validation part of the ROUGH dataset (Section IV-A) following the procedure described in Section III-D. As the evaluation metrics, we use terrain prediction accuracy, Table I. The $L2$-norms are computed to estimate geometrical, $\Delta\mathcal{H}_{g}$ (9), and terrain, $\Delta\mathcal{H}_{t}$ (10), heightmap prediction accuracy covering both geometric- and semantic- oriented tasks. The results in the two right-most columns of the Table I contain the terrain prediction accuracy metrics. The lidar input-based model (VoxelNet) performs better than the camera input-based LSS. However, the benefit of the sensor fusion on the terrain accuracy prediction is not obvious as the metrics $\Delta\mathcal{H}_{g}$ and $\Delta\mathcal{H}_{t}$ for both models (VoxelNet and BEVFusion) are nearly equal.
Figure 14: The supporting terrain $\mathcal{H}_{t}$ prediction with the LSS [20] (first row), VoxelNet [21] (second row), and BEVFusion [22] (third row) models. The predicted supporting terrain $\mathcal{H}_{t}$ is projected onto camera images (as a point cloud) and visualized in the top-down view (on the right).
@@
The qualitative results of the terrain properties prediction are visualized in Figure 14. It can be seen that although the LSS model provides a good estimate of the terrain shape (and thus the predicted trajectory matches the ground truth quite closely), it struggles to predict obstacles with sharp edges (tree trunks). The models that have a lidar input (VoxelNet and BEVFusion) provide more accurate terrain predictions. The BEVFusion prediction is also visualized in 3D in Figure 1 with the lidar point cloud input.
TABLE I: Trajectory and terrain estimation accuracy.
input method terrain encoder $\tau$ pred. $\Delta\mathbf{x}$ [$\mathrm{m}$] $\Delta\mathbf{R}$ [$\deg$] $\Delta\mathcal{H}_{g}$ [$\mathrm{m}$] $\Delta\mathcal{H}_{t}$ [$\mathrm{m}$]
[…]
RGB data-driven LSS [20] TrajLSTM [19] 0.128 3.949
[…]
point cloud data-driven VoxelNet [21] TrajLSTM [19] 0.129 3.369
[…]
RGB + point cloud data-driven BEVFusion [22] TrajLSTM [19] 0.125 3.067
@@ IV-C Physics-based vs Data-driven Baseline
${}\Delta\mathbf{R}=\frac{1}{T}\sum_{t=0}^{T}\arccos\frac{\mathrm{tr}({\mathbf{R}^{\top}_{t}\mathbf{R}^{*}_{t}})-1}{2}$ (12)
@@ IV-D Physics Engine Computational Efficiency
(a) CPU: AMD Ryzen 7 4800H, GPU: GeForce GTX 1660 Ti Mobile, $N_{traj}=512$
(b) GPU: Tesla V100-SXM2-32GB, $N_{traj}=2048$
@@
The Figure 15 shows the runtime of the differentiable physics engine depending on the predicted trajectories time horizon for the two solvers’ implementation methods and different hardware configurations. The Figure 15 (a) shows the runtime on the standard laptop hardware (CPU and GPU), while the Figure 15 (b) shows the runtime on the high-performance GPU. It can be noticed that the runtime grows linearly with the time horizon in all cases and the laptop GPU implementation is around 5 times faster than the CPU one. The Neural ODE solver [44] is slightly more efficient than the auto-differentiation one. For example, on a high-performance GPU (Tesla V100-SXM2-32GB), the prediction of 2048 trajectories ($6~{}[$\sec$]$-long) takes around $1~{}[$\sec$]$. The short runtime of the differentiable physics engine allows for its efficient usage in the end-to-end learning pipeline (Section III-D) and for real-time trajectory shooting in navigation scenarios (Section IV-E).
@@ IV-E Navigation in Unstructured Environments
Figure 17: Autonomous navigation in the forest environment with MonoForce. The robot follows a set of waypoints and at the same time avoids obstacles (trees, bushes, rocks, etc.). First row (Control setup): MonoForce prediction, based only on the onboard camera, includes terrain shape and a set of trajectories for different control commands sampled through a simplified MPPI technique. The colors of a possible trajectory correspond to the cost (red is the most expensive). The trajectory with green arrows is the one with the smallest total cost, (15). Second row (Qualitative results): Predicted supporting terrain and robot trajectory given camera images. The predictions and frontal camera images correspond to different time moments of the navigation experiment in the forest environment (Figure 16).
Figure 18: MonoForce prediction projected to robot’s camera images. The prediction includes supporting terrain $H_{t}$; the colors correspond to the friction coefficient $\mu$ (violet is low, red is high). A set of 64 trajectories for different control commands is visualized in white and black (low and high trajectory cost (13) respectively). The black frames denote the camera’s field of view.
Thanks to integration of the extereoceptive sensor to Terrain Encoder and GPU parallelization of the predicted trajectories (Section III-B), it is possible to use the MonoForce model for autonomous navigation in unstructured outdoor environments. In the navigation experiments, the robot is given a set of distant waypoints to follow. It is localized using the ICP SLAM method [47]. The MonoForce model runs onboard the robot’s hardware and predicts terrain properties in front of the robot and a set of possible trajectories for different control commands. In the case of a tracked robot platform (Figure 12), the controls are the commanded velocities of individual tracks. A sample of predicted trajectories for given controls is visualized in Figure 18. To reach a goal safely, the robot chooses the trajectory with the smallest cost and distance to the next waypoint, Figure 17 Thanks to the MonoForce ability to predict robot-terrain interaction forces for a trajectory pose, it is possible to estimate the cost of each trajectory. We decide to calculate the cost as a values variance of reaction forces acting on the robot along a trajectory $\tau$:

### v4

@@ III-B Differentiable Neuro-Symbolic Physics Engine
When the robot is moving over a terrain, two types of external forces are acting on the point cloud $\mathcal{P}$ representing its model: (i) gravitational forces and (ii) robot-terrain interaction forces. The former is defined as $m_{i}\mathbf{g}=[0,0,-m_{i}g]^{\top}$ and acts on all the points of the robot at all times, while the latter is the result of complex physical interactions that are not easy to model explicitly and act only on the points of the robot that are in contact with the terrain. There are two types of robot-terrain interaction forces: (i) normal terrain force that prevents the penetration of the terrain by the robot points, (ii) tangential friction force that generates forward acceleration when the tracks are moving, and prevents side slippage of the robot, see Fig. 6 for a simplified 2D sketch.
@@
Normal reaction forces: One extreme option is to predict the 3D force vectors $\mathbf{F}_{i}$ directly by a neural network, but we decided to enforce additional prior assumptions to reduce the risk of overfitting. These prior assumptions are based on common intuition from the contact dynamics of flexible objects. In particular, we assume that the magnitude of the force that the terrain exerts on the robot’s point $\mathbf{p}_{i}\in\mathcal{P}$ increases proportionally to the deformation of the terrain. Consequently, the network does not directly predict the force, but rather predicts the height of the terrain $h\in\mathcal{H}_{t}$ at which the force begins to act on the robot body and the stiffness of the terrain $e\in\mathcal{K}$. We understand the quantity $e$ as an equivalent of the spring constant from Hooke’s spring model, Figure 6. Given the stiffness of the terrain and the point of the robot that penetrated the terrain by ${\Delta}h$, the reaction force is calculated as $e\cdot{\Delta}h$.
@@
Tangential friction forces: Our tracked robot navigates by moving the main tracks and 4 flippers (auxiliary tracks). The flipper motion is purely kinematic in our model. This means that in a given time instant, their pose is uniquely determined by a $4$-dimensional vector of their rotations, and they are treated as a rigid part of the robot. The motion of the main tracks is transformed into forces tangential to the terrain. The friction force delivers forward acceleration of the robot when robot tracks (either on flippers or on main tracks) are moving. At the same time, it prevents the robot from sliding sideways. When a robot point $\mathbf{p}_{i}$, which belongs to a track, is in contact with terrain with friction coefficient $\mu\in\mathcal{M}$, the resulting friction force at a contact point is computed as follows, [38]:
@@
Implementation of the Differentiable ODE Solver: We implement the robot-terrain interaction ODE (5) in PyTorch [47]. The Neural ODE framework [48] is used to solve the system of ODEs. For efficiency reasons, we utilize the Euler integrator for the ODE integration. The differentiable ODE solver [48] estimates the gradient through the implicit function theorem. Additionally, we implement the ODE (5) solver that estimates gradient through auto-differentiation [47], i.e. it builds and retains the full computational graph of the feedforward integration.
@@ III-C Data-driven Trajectory Prediction
Given an initial robot’s state $\mathbf{x}_{0}$ and a sequence of control inputs for a time horizon $T$, $\mathbf{u}_{t},t\in\{0\dots T\}$, the TrajLSTM model provides a sequence of states at control command time moments, $\mathbf{x}_{t},t\in\{0\dots T\}$. As in outdoor scenarios the robot commonly traverses uneven terrain, we additionally include the terrain shape input to the model in the form of heightmap $\mathcal{H}=\mathcal{H}_{0}$ estimated at initial time moment $t=0$. Each timestep’s control input $\mathbf{u}_{i}$ is concatenated with the shared spatial features $\mathbf{x}_{i}$, as shown in Fig. 8. The combined features are passed through dense layers to prepare for temporal processing. The LSTM unit [32] processes the sequence of features (one for each timestep). As in our experiments, the time horizon for trajectory prediction is reasonably small, $T=5[$\sec$]$, and the robot’s trajectories lie within the heightmap area, we use the shared heightmap input for all the LSTM units of the network. So the heightmap is processed through the convolutional layers once and flattened, producing a fixed-size spatial feature vector. This design choice (of not processing the heightmaps at different time moments) is also motivated by computational efficiency reason. At each moment $t$, this heightmap vector is concatenated with the fused spatial-control features and processed by an LSTM unit. The LSTM unit output for each timestep $t$ is passed through a fully connected (dense) layer to produce the next state $\mathbf{x}_{t+1}$. The sequence of states form the predicted trajectory, $\{\mathbf{x}_{0},\dots\mathbf{x}_{T}\}$.
@@ III-D End-to-end Learning
Since the architecture Fig. 2 is end-to-end differentiable, we can directly learn to predict all intermediate outputs just using trajectory loss (10). An example of terrain learning with the trajectory loss is visualized in Fig. 7. To make the training more efficient and the learned model explainable, we employ the geometrical loss (11) and terrain loss (12) as regularization terms. stat
@@
The Fig. 10 show the prediction examples of the FusionForce model in diverse outdoor environments. From the example on the left, we can see that the model correctly predicts the robot’s trajectory and the terrain shape suppressing traversable vegetation, while the rigid obstacles (wall and tree logs) are correctly detected. The example on the right demonstrates the model’s ability to predict the robot’s trajectory ($10~{}[$\sec$]$-long horizon) with reasonable accuracy and to detect the rigid obstacles (stones) on the terrain. It could also be noticed that the surfaces that provide the robot good traction (paved and gravel roads) are marked with a higher friction value, while for the objects that might not give good contact with the robot’s tracks (walls and tree logs) the friction value is lower.
@@
We argue that the friction estimates are approximate and an interesting research direction could be comparing them with real-world measurements or with the values provided by a high-fidelity physics engine (e.g. AGX Dynamics [42]). However, one of the benefits of our differentiable approach is that the model does not require ground-truth friction values for training. The predicted heightmap’s size is $12.8\times 12.8$\mathrm{m}$^{2}$ and the grid resolution is $0.1$\mathrm{m}$$. It has an upper bound of $1~{}[$\mathrm{m}$]$ and a lower bound of $-1~{}[$\mathrm{m}$]$. This constraint was introduced based on the robot’s size and taking into account hanging objects (tree branches) that should not be considered as obstacles (Fig. 15). Additionally, the terrain is predicted in the gravity-aligned frame. That is made possible thanks to the inclusion of camera intrinsics and extrinsics as input to the model, Fig. 5. It also allows correctly modeling the robot-terrain interaction forces (and thus modeling the robot’s trajectory accurately) for the scenarios with non-flat terrain, for example, going uphill or downhill. This will not be possible if only camera images are used as input.
IV Results
TABLE I: Trajectory and terrain estimation accuracy. RGB = input is onboard camera image, PCL = input is onboard lidar pointcloud only, ”hybrid” = architecture combining classical layers with the proposed neural symbolic physics engine, ”data-driven” = architecture employing classical layers (feedforward and recurrent) only.
input method terrain encoder $\tau$ pred. $\Delta\mathbf{x}$ [$\mathrm{m}$] $\Delta\mathbf{R}$ [$\mathrm{rad}$] $\Delta\mathcal{H}_{g}$ [$\mathrm{m}$] $\Delta\mathcal{H}_{t}$ [$\mathrm{m}$]
[…]
RGB data-driven LSS TrajLSTM 0.128 0.070
PCL data-driven VoxelNet TrajLSTM 0.129 0.054
RGB+PCL data-driven BEVFusion TrajLSTM 0.125 0.053
Figure 10: Qualitative examples. Left: The robot is moving through a narrow passage between a wall and tree logs. Right: The robot is moving on a gravel road with rocks on the sides. It starts its motion from the position marked with a coordinate frame and the trajectory is predicted for $10~{}[$\sec$]$ using real control commands. The camera images are taken from the robot’s initial position (top row). The visualization includes predicted supporting terrain $\mathcal{H}_{t}$ (second row). It is additionally shown in 3D and colored with predicted friction values (third row).
IV-A Datasets
@@
Despite the amount and annotation quality of the data provided in the RELLIS-3D sequences, it lacks examples of a robot moving over hills, obstacles, and traversing high grass. To fill this gap we manually collected several hours of driving with the flipper-based skid-steering robot in challenging terrain. This dataset contains traversals on (or through) flexible natural obstacles like tall grass, hanging tree branches, mud, dense undergrowth, see Figure 11 for an example. Recorded traversal contains dynamically changing flipper positions to cover the variability of underlying physics problem. We refine RELIS-3D-pretrained FusionForce on these new data.
[…]
@@ IV-B Sensor Fusion
We train and evaluate the proposed FusionForce model with the four state-of-the-art baseline architectures. The first three considered architectures consist of LSS [27], VoxelNet [51], and BEVFusion [52] used as terrain encoders, followed by a data-driven trajectory predictor LSTM [36]. We refer to these architectures as data-driven. The last considered baseline is our own previous work called MonoForce [19] that uses $\nabla$-physics neuro symbolic layer for trajectory prediction. We refer to this architecture as hybrid. All networks are trained using the losses described in Section III-D on the same training and evaluation data.
All approaches are compared in terms of two different metrics: (i) accuracy of trajectory prediction and (ii) accuracy of terrain properties prediction. The former represents accuracy computed on real trajectories, which is inherently limited only to safe traversals/terrains, while the latter mostly represents accuracy on dangerous terrains. Consequently, both metrics are equally important. Table I summarizes the results.
The trajectory prediction accuracy is reported as translational and rotational error computed as follows:
${}\Delta\mathbf{R}=\frac{1}{T}\sum_{t=0}^{T}\arccos\frac{\mathrm{tr}({\mathbf{R}^{\top}_{t}\mathbf{R}^{*}_{t}})-1}{2},$ (14)
where $\{\mathbf{x}_{t}$ is predicted translation, $\mathbf{R}_{t}\}$ rotation and $\{\mathbf{x}_{t}^{*},\mathbf{R}_{t}^{*}\}$, $t\in\{0\dots T\}$ is the ground truth translation and rotation. The ground truth poses were recorded using the SLAM method introduced in [53].
Similarly, the terrain prediction accuracy is evaluated as the $L2$-error of geometrical $\Delta\mathcal{H}_{g}$ (11), and terrain $\Delta\mathcal{H}_{t}$ (12) heightmaps on evaluation data. These metrics are shown in the two right columns of the Table I. We observe that FusionForce achieves, independently of the employed input modality, consistently smaller terrain prediction and trajectory estimation errors compared to the MonoForce baseline. The improvement in terrain accuracy suggests a better generalization on dangerous (out-of-distribution) terrains.
@@
The qualitative results of the prediction of terrain properties are visualized in Fig. 4. Although the camera-only model provides a good estimate of the terrain shape (and thus the predicted trajectory matches the ground truth quite closely), it struggles to predict obstacles with sharp edges such as tree trunks. Using the lidar pointcloud as input provides more accurate terrain predictions.
@@ IV-C Physics-based vs Data-driven Baseline
[…]
@@ IV-D Physics Engine Computational Efficiency
Figure 15: Autonomous navigation in the forest environment. The robot follows a set of waypoints and at the same time avoids obstacles (trees, bushes, rocks, etc.). First row (Control setup): Trajectory predictions, based only on the onboard camera, for different control commands are sampled through a simplified MPPI technique. The colors of a possible trajectory correspond to the cost (red is the most expensive), and the yellow cross denotes the position of the closest waypoint. The trajectory with green arrows is the one with the smallest total cost, (17). Second row (Qualitative results): Predicted supporting terrain and robot trajectory given camera images. The predictions and frontal camera images correspond to different time moments of the navigation experiment in the forest environment (Fig. 14).
The Fig. 13 shows the runtime of the differentiable physics engine depending on the predicted trajectories time horizon for the two solvers’ implementation methods and different hardware configurations. It can be noticed that the runtime grows linearly with the time horizon in all cases and the V100 GPU implementation is around 10 times faster than the CPU one. The Neural ODE solver [48] is slightly more efficient than the auto-differentiation one. For example, on a high-performance GPU (Tesla V100-SXM2-32GB), the prediction of 2048 trajectories ($6~{}[$\sec$]$-long) takes around $1~{}[$\sec$]$. The short runtime of the differentiable physics engine allows for its efficient usage in the end-to-end learning pipeline and for real-time trajectory shooting in navigation scenarios.
IV-E Autonomous Navigation in Unstructured Environments
We demonstrate that the proposed model can be easily used for autonomous navigation in unstructured outdoor environments. In the navigation experiments, the robot is given a set of distant waypoints to follow and it is localized using the ICP SLAM method [53]. The model runs onboard the robot’s hardware and predicts terrain properties in front of the robot and a set of possible trajectories for different control commands. In the case of a tracked robot platform (Fig. 9), the controls are the commanded velocities of individual tracks. Sample of predicted trajectories for given controls is visualized in Fig. 15. To reach a goal safely, the robot chooses the trajectory with the smallest cost and distance to the next waypoint.
Thanks to the ability of the FusionForce to predict robot-terrain interaction forces for a trajectory pose, it is possible to estimate the cost of each trajectory. We decide to calculate the cost as a values variance of reaction forces acting on the robot along a trajectory $\tau$: